OpenStudy (anonymous):

EF is a median of trapezoid ABCD. What is the value of x? (Let me draw the trapezoid first)

OpenStudy (anonymous):

 this is the trapezoid

OpenStudy (anonymous):

I NEED HELP @aum @Ibbutibbu. @TheSaint905621

OpenStudy (ibbutibbu.):

ok lets see...

OpenStudy (ibbutibbu.):

wow this is hard

OpenStudy (ibbutibbu.):

ok so one side has to be less than ten and the other has to be more than ten

OpenStudy (ibbutibbu.):

there is no other information in the question..?

OpenStudy (anonymous):

no

OpenStudy (ibbutibbu.):

hmm...

OpenStudy (ibbutibbu.):

i think... OK lets try substituting random numbers for x. lets strart with zero

OpenStudy (anonymous):

ok

OpenStudy (ibbutibbu.):

it cannot be zero because the top line has to be positive.

OpenStudy (ibbutibbu.):

so lets try one

OpenStudy (anonymous):

ok

OpenStudy (ibbutibbu.):

the only way you can fin the area is to find the height

OpenStudy (ibbutibbu.):

and with the area you could probably figure out the x value

OpenStudy (ibbutibbu.):

hmm. @satellite73 , can you help?

OpenStudy (anonymous):

yeah and maybe @NaeNaeHoowaahh or @TheSaint905621

OpenStudy (anonymous):

the median is always the difference of the two bases

OpenStudy (ibbutibbu.):

your GOOD

OpenStudy (anonymous):

thankyou

OpenStudy (aum):

I think EF = (AB + DC) / 2

OpenStudy (anonymous):

Hey maybe @LearningIsAwesome can help too

OpenStudy (anonymous):

I'm sorry I'm really bad a Geometry, I'm good with equations and basic statistics and some polynomials, other than that in the math thing I'm not the best on...

OpenStudy (anonymous):

maybe @lxl-rhea-lxl can help

OpenStudy (aum):

EF = (AB + DC) / 2 = (3x - 1 + 7x + 1) / 2 = (10x) / 2 = 5x 10 = 5x x = 2

OpenStudy (anonymous):

@mathstudent55 can help I think! Love his explanations

OpenStudy (ibbutibbu.):

hey lets give each other medals

OpenStudy (anonymous):

That's against the rules...

OpenStudy (ibbutibbu.):

oh sry

OpenStudy (ibbutibbu.):

i had no clue

OpenStudy (anonymous):

Sorry, I had to report it.

OpenStudy (anonymous):

u know what let's try a new question

OpenStudy (ibbutibbu.):

its all right :)

OpenStudy (aum):

Aha! Found a confirmation for my hunch: http://www.regentsprep.org/Regents/math/geometry/GP9/LTrapezoid.htm